# Chirality Amplification and Deracemization in Liquid Crystals

### v1

@@ Results
The helical twisting power efficacy of the same dopant in different nematogens may be interpreted in terms of the distance in temperature to the chiral transition of that nematogen, yielding different values of $\chi$, and differences in the binding free energy, $g$, and the intrinsic helical wave number, $q_{0}$. Changes in the binding free energy between different dopants result in different helical twisting powers dependent on the nematogen. Novel in this expression for the helical twisting power is the pretransitional enhancement factor $(1-\chi)^{-1}$ not accounted for in other approaches [18].
It turns out to be nontrivial to evaluate the temperature dependence of the helical twisting power, as $\chi$, $g$ and $q_{0}$ all depend on the temperature. For our model, $q_{0}=K/2J$ is often thought to not depend on temperature [13] but this presumes that energies $K$ and $J$ are not free energies. Since the interaction energy is based on a coarse-grained model, $K$ and $J$ must in fact be interpreted as free energies in which case both are in principle temperature dependent. If we ignore this for the sake of argument, and presume $K$ and $J$ to be a constant of temperature, then we deduce from Eq. 4 that $\beta$ should decrease with temperature under most (but not all) conditions, because higher temperatures imply larger distance to the critical deracemization temperature and smaller values of $\chi$. This is often seen experimentally but not always [19]. However, if the binding of the dopant to the nematogen is entropy- rather than enthalpy-driven, and $g/k_{B}T$ increases with temperature, then it may actually increase with temperature. This could explain recent experimental observations [20]. We refer to the Supporting Information for a more detailed discussion of the factors that influence the temperature dependence of the helical twisting power $\beta$.
@@
To explore cooperative chiral ordering and spontaneous deracemization, we study by means of Monte Carlo simulation a switchable chiral Lebwohl-Lasher model as described by the following Hamiltonian:
@@
where the sum is over neighboring pairs $\langle i,j\rangle$ of lattice sites, $J$ and $K$ have the same meaning as before, as does the biasing field $g$, and $\hat{r}_{ij}$ denotes the unit vector connecting these two sites. Associated with each site are two order parameters: the unit orientation vector $\hat{u}_{i}$ and the chirality $\eta_{i}=\pm 1$ describing the left or right handedness of the mesogens. The first term in Eq. 5 represents the usual Lebwohl-Lasher potential [21, 22], and the second term is a chiral interaction term recently explored in a different context by Elsasser and Kuhnhold [23] without the local switchable chirality that is present in our model. The coupling term $\eta_{i}+\eta_{j}$ promotes twist between neighboring “spins” for pairs of like chirality $\eta_{i}=\eta_{j}$ but not between neighboring spins of opposite chirality $\eta_{i}\neq\eta_{j}$. More simulation details are available in the methods section. Simulations were performed on the National Center for Supercomputing Applications’ Delta cluster through the NSF ACCESS program [24]. A copy of the simulation code is available at [25].
@@
We simulate the model using an $N\times N\times N$ cubic lattice with $N=128$. To explore the phase diagram, the chiral interaction strength, $K/J$, was varied from $K/J=0.0$ to $K/J=2.0$ with a total of $13$ values. For each value of $K/J$, the temperature was varied from $T/T_{IN}=2.0$ to $T/T_{IN}=0.067$ with $2\times 10^{6}$ Monte Carlo steps at each of the $30$ temperature values. For each set of parameters, we used the last $6.25\%$ of configurations to calculate specific heat and chiral susceptibility, with configurations separated by 100 Monte Carlo steps per spin. The specific heat was calculated from the energy fluctuations as $C_{V}=\left(\langle E^{2}\rangle-\langle E\rangle^{2}\right)/k_{\mathrm{B}}T^{2}$. The chiral susceptibility is calculated from the fluctuations in the enantiomeric excess $\eta$, $\chi_{\eta}=N^{3}\left(\langle\eta^{2}\rangle-\langle\eta\rangle^{2}\right)/k_{\mathrm{B}}T$. The enantiomeric excess was calculated as the average chirality per lattice site, in direct analogy to the average spin state in the Ising model.
@@
The phase diagram obtained from our simulations is shown in Fig.3 (a), showing peaks in the heat capacity that indicate transitions between racemic isotropic, racemic nematic and deracemized cholesteric phases, in close agreement with predictions of our Maier-Saupe theory of Fig. 1. We find a racemic isotropic phase at high temperature, a racemic nematic phase at low temperature and sufficiently weak chiral interaction $K/J$, and a spontaneously deracemized cholesteric phase at low temperature and high $K/J$. The intersection of the phase boundaries suggests the existence of a tricritical point, also predicted by our Maier-Saupe theory. We in addition plot the chiral susceptibility in Fig. 3 (b), which can only show the phase boundary of the spontaneously deracemized cholesteric phase. It confirms that near the transition, chiral fluctuations become large and hence also the chiral susceptibility.
@@
Even not all that close to the transition, the increased susceptibility can be quite significant as Fig. 2 (b) confirms. In the figure, the enantiomeric excess $\eta$ is shown as a function of the dimensionless chiral bias field for different temperatures but fixed chiral interaction strength $K/J=0.333$. In our Monte Carlo simulations, we varied $g/k_{\mathrm{B}}T$ from $0.0$ to $0.25$, $K/J$ from 00 to $0.5$, and $T/T_{IN}$ from $0.668$ to $0.267$, where we annealed $2\times 10^{6}$ Monte Carlo steps at each temperature. In agreement with our theory, we find that the chiral susceptibility goes up with decreasing temperature and with increasing chiral interaction strength $K/J$, as the Fig. S1 shows. According to our theory, both $T$ and $K/J$ enter because $\eta$ not only depends on $g/k_{\mathrm{B}}T$ but also on the value of $\chi$ that itself is a function of $T/T_{\mathrm{IN}}$ and $K/J$.
@@
What Maier-Saupe theory cannot show, and our simulations can, is the spatial structure of the phases and their coarsening behavior. Fig. 4 (a) and (b) show $xz$ cross-sections of a system with size $N=128$, with open boundary conditions on all sides. Both simulations were annealed to a temperature of $T/T_{IN}=0.067$ starting from $T/T_{IN}=1.0$, in $30$ steps of $2\times 10^{6}$ Monte Carlo steps each. In the figures, right-handed ${\eta_{i}}=+1$ spins are shown in red and left-handed ${\eta_{i}}=-1$ spins are shown in blue. Fig. 4(a) shows the racemic nematic phase at low value of $K/J=0.167$, with short range chiral correlations, but no long range order and no induced cholesteric twist. Fig. 4(b) is at a higher value of $K/J=0.5$, showing the deracemized phase with layers of cholesteric of approximately two pitch lengths each. A more detailed image can be found in Fig. S3 demonstrating that under these conditions the interface between left- and right-handed deracemized cholesteric layer is indeed very sharp. The cholesteric domains form with the pitch axis oriented along a body diagonal of the cubic lattice, see also Fig. 4 (c). This orientation of the pitch axis is allowed by open boundary conditions and is lower in energy than a pitch oriented along the $x$, $y$ or $z$ axes of the cubic simulation volume.
To investigate coarsening behavior in the deracemized phase, we carried out a simulation of a larger system with $N=480$ and open boundary conditions, for a fixed chiral interaction ${K/J}=0.5$ and a temperature $T/T_{IN}=0.067$ that falls in the deracemized cholesteric region of the phase diagram. We followed the coarsening over $2\times 10^{6}$ Monte Carlo steps per spin. Our findings are shown in Fig. 4 (c) and (d), where layers with ${\eta_{i}}=1$ are again shown in red and with ${\eta_{i}}=-1$ in blue. Layers with alternating chiral sense form within the first $1\times 10^{5}$ time steps after starting from an initial random configuration as seen in Fig. S5 (a). These widen with time as can be seen in more detail a in the cross sections shown in Fig. S2 and Fig. S5. Initially, multiple domains form with layer normals aligned along different body diagonal axes. Over time, those domains grow with time, the average layer width increases, and the number of layers drop. Layers annihilate via two mechanisms: (i) via motion and annihilation of edge dislocations in the layer structure, and (ii) via layer annihilation at corners of the cubic box facilitated by the open boundary conditions. The average layer thickness $L$ we find to increase continuously as $L\sim{t^{\alpha}}$ with scaling exponent $\alpha={0.379}\pm{0.016}$ as shown in Fig. 4 (d). Because layer formation and coarsening are driven by an intrinsically anisotropic surface tension, we do not expect this scaling exponent to be universally valid [26, 27].
As Fig. 4 (c) shows, the domain structure is actually quite complex, each with layer normals oriented along different body diagonals. Layers formed in early stages of pattern formation have thicknesses less than the cholesteric pitch, increasing to several pitch lengths during coarsening. See Fig. S3 and Fig. S6 b). Layers less wide than about a pitch do not have a simple, single twisted cholesteric structure but exhibit a double-twist structure. Layers annihilate by shrinking to zero thickness at a corner or edge of the simulation box, or via motion and annihilation of edge dislocations. The system remains in a poly-domain structure after $2\times{10^{6}}$ Monte Carlo steps per spin. The dynamics of micro-structural evolution is shown in Supplementary Video S1 and in Fig. S5. Cross-sectional images are shown in Fig. S2.
Finally, for values of $K/J=0.833$ and larger, a more complex, disordered double twist geometry without a uniform twist axis emerge. The defect-rich micro-structures suggest the formation of blue phases, which have not previously been observed in studies of the chiral Lebwohl-Lasher model. We plan to investigate the details of the blue phase in the uniform chirality Lebwohl-Lasher model in future work.
Discussion
Spontaneous deracemization may occur by a variety of mechanisms [28], and is not restricted to racemic mixtures of chiral molecules. In fact, achiral mesogens can form chiral phases even in the absence of intramolecular degrees of freedom. For instance, bent-core molecules self-assemble into smectic phases with a chiral arrangement of the layer normals, the molecular dipole orientations, and molecular tilts [29, 30]. Also, twist-bend nematic phases form without reliance of intramolecular degrees of freedom [31, 32, 33, 34, 35]. Intramolecular degrees of freedom can also drive deracemization [15]. While molecular chirality typically arises due to the presence of a chiral center, an achiral molecule with degenerate right- and left-handed twist conformations can be described as having axial chirality [8]. Here two mirror-image chiral conformations differ only by the sign of axial twist in the molecular core.
In the model we have presented, our focus is on mesophases of molecules with axial chirality where the energy barrier between mirror-image conformations is sufficiently small relative to the thermal energy, enabling switching between chiral states. Our model also assumes that interactions of neighboring molecules promote agreement in their axial chiral state and predicts segregation into chiral domains. Wilson et al.  have clearly demonstrated this mechanism via atomistic molecular simulations of flexible bent-core mesogens [12]. In related work there is evidence that intramolecular degrees of freedom in bent-core molecules enables them to form chiral conformations with resulting impact on their phase behavior as well as their chiral effect as dopants on other mesophase materials [36].
In summary, by making use of a relatively simple Maier-Saupe theory that allows for a thermodynamic equilibrium to exist between conformers of opposite handedness, we are able to explain chirality amplification in nematic liquid crystals. According to our model, chirality amplification is caused by an underlying tendency of enantiomers with opposite handedness to demix, which makes them susceptible to a chiral field even when the second order demixing transition is obscured by a transition to another liquid or solid crystal phase. In follow-up work, we intend to study the phase behavior of our model in much more detail and focus on potential dynamical effects associated with the chiral hysteresis predicted by our theory, in analogy to the hysteresis observed in the magnetization of ferromagnets.
Methods
Theory
Following van der Meer et al. [13], we find from a mean-field type of analysis of the impact of a chiral Lebwohl-Lasher type of interaction potential between model mesogens to lead to expressions for the self-consistent fields experienced by species $\alpha=\pm$ due to the presence of species $\beta=\pm$ to read
The chiral coupling constants $K_{\alpha,\beta}$ are $K_{+,+}=K$ between the $+$ species, $K_{-,-}=-K$ between the -- species and $K_{+,-}=K_{-,+}=0$ between the species with opposite handedness. In Eq. 6, we have absorbed unimportant numerical constants in the definitions of the interaction strengths $J$ and $K$ as well as the cholesteric wave number $q$ that in our description is made dimensionless by multiplication with a microscopic length scale [13]. The free energy of interaction follows from summing all self-consistent fields experienced by species $\alpha$ multiplied with the appropriate fraction $f_{\alpha}$, and accounting for double counting. Adding the Gibbs entropy for the orientational distribution and that for the distribution over the species $\alpha$, as well as the contribution of the biasing field $g$, then produces, after some rearranging, Eq. 1.
Simulation Methods
To explore cooperative chiral ordering and spontaneous deracemization by means of Monte Carlo computer simulation, we study a switchable chiral 3D Lebwohl-Lasher model based on the Hamiltonian of Eq. 5. We perform Monte Carlo simulations of the model via the Metropolis algorithm, where at each Monte Carlo step we separately vary the local orientation $\hat{u}_{i}$ and chiral state parameter $\eta_{i}$. We use a novel method for trial spin reorientation which realizes a smoother reorientation of the spin, rather than the Barker-Watts technique [37] used in [22]. To perform a random walk on a unit sphere, we calculate two vectors orthogonal to the orientation vector $\hat{u}_{i}$ and tangent to the unit sphere, then reorient the spin to a random position on the edge of a disk of radius $d$, and normalize the orientation vector to unit length. The disk size $d$ represents the step size and is adjusted to keep the Monte Carlo acceptance rate in the range of 0.4 to 0.6 during equilibration at each temperature, and after that is held fixed. We initialize each simulation in the racemic isotropic phase.
Lattice-based Monte-Carlo simulations are a computationally “embarrassingly parallelizable” problem [38]. Our Monte Carlo code is highly optimized to run on graphics processing units (GPUs) using a combination of Fortran and OpenACC for the bulk of the computation, and Python for simulation management, data processing, analysis, and visualization. Simulations were performed on the National Center for Supercomputing Applications’ Delta cluster through the NSF ACCESS program [24]. A copy of the simulation code is available at doi.org/10.5281/zenodo.14781846. [25].

### v2

@@ Results
The helical twisting power efficacy of the same dopant in different nematogens may be interpreted in terms of the distance in temperature to the chiral transition of that nematogen, yielding different values of $\chi$, and differences in the binding free energy, $g$, as well as the intrinsic helical wave number, $q_{0}$. Changes in the binding free energy between different dopants result in different helical twisting powers dependent on the nematogen. Novel in our expression for the helical twisting power Eq. 5 is the pretransitional enhancement factor $(1-\chi)^{-1}$ not accounted for in other approaches [18]. See also the SI.
The temperature dependence of the helical twisting power that we predict is not universal, as $\chi$, $g$ and $q_{0}$ all depend on the temperature. For our model, $q_{0}=K/2J$ is often thought to not depend on temperature [46] but this presumes that $K$ and $J$ are energies. However, because the interaction energy is that of a coarse-grained model in which degrees of freedom have been “integrated out”, $K$ and $J$ must in fact be interpreted as free energies that are temperature dependent. If we ignore this and presume $K$ and $J$ to be a constant of temperature, we deduce from Eq. 5 that $\beta$ should decrease with temperature under most conditions because higher temperatures imply larger distance to the critical chiral segregation temperature and smaller values of $\chi$. This is indeed usually seen experimentally, but not always [26]. However, if the binding of the dopant to the nematogen is entropy- rather than enthalpy-dominated, and $g/k_{B}T$ increases with temperature, then $\beta$ may actually increase with temperature. This could explain recent observations [51]. We refer to the SI for a more detailed discussion of the factors that influence the temperature dependence of the helical twisting power $\beta$.
@@
To explore cooperative chiral ordering and spontaneous chiral segregation, we study by means of Monte Carlo simulation a switchable chiral Lebwohl-Lasher model as described by the following Hamiltonian:
@@
where the sum is over neighboring pairs $\langle i,j\rangle$ of lattice sites, $J$ and $K$ have the same meaning as before, as does the biasing field $g$, and $\hat{r}_{ij}$ denotes the unit vector connecting these two sites. Associated with each site are two order parameters: the unit orientation vector $\hat{u}_{i}$ and the chirality $\eta_{i}=\pm 1$ describing the left or right handedness of the mesogens. The first term in Eq. 6 represents the usual Lebwohl-Lasher potential [27, 40], and the second term is a chiral interaction term which was recently explored in a different context by Elsasser and Kuhnhold [14] without the local switchable chirality part of our model. Our model incorporates a coupling term $\eta_{i}+\eta_{j}$, promotingtwist between neighboring “spins” for pairs of like chirality $\eta_{i}=\eta_{j}$ but not between neighboring spins of opposite chirality $\eta_{i}\neq\eta_{j}$. More simulation details are available in the Methods section. Simulations were performed on the National Center for Supercomputing Applications’ Delta cluster through the NSF ACCESS program [5]. A copy of the simulation code is available at [8].
@@
We simulate the model on an $N\times N\times N$ cubic lattice with $N=128$. To explore the phase diagram, the chiral interaction strength, $K/J$, was varied from $K/J=0.0$ to $K/J=2.0$ with a total of $13$ values. For each value of $K/J$, the temperature was varied from $T/T_{IN}=2.0$ to $T/T_{IN}=0.067$ with $2\times 10^{6}$ Monte Carlo steps at each of the $30$ temperature values. For each set of parameters, we used the last $6.25\%$ of configurations to calculate specific heat and chiral susceptibility, with configurations separated by 100 Monte Carlo steps per spin. The specific heat was calculated from the energy fluctuations as $C_{V}=\left(\langle E^{2}\rangle-\langle E\rangle^{2}\right)/k_{\mathrm{B}}T^{2}$. The chiral susceptibility is calculated from the fluctuations in the enantiomeric excess $\eta$, $\chi_{\eta}=N^{3}\left(\langle\eta^{2}\rangle-\langle\eta\rangle^{2}\right)/k_{\mathrm{B}}T$. The enantiomeric excess was calculated as the average chirality per lattice site, in direct analogy to the average spin state in the Ising model.
@@
The phase diagram obtained from our simulations is shown in Fig. 3 (a), showing peaks in the heat capacity that indicate transitions between racemic isotropic, racemic nematic and segregated cholesteric phases with opposite handedness and enantiomeric excess, in close agreement with predictions of our Maier-Saupe theory of Fig. 1. We find a racemic isotropic phase at high temperature, a racemic nematic phase at low temperature and sufficiently weak chiral interaction $K/J$, and spontaneously segregated cholesteric domains at low temperature and high $K/J$. The intersection of the phase boundaries points at the existence of a critical end point, also predicted by our Maier-Saupe theory. In addition, we plot the chiral susceptibility in Fig. 3 (b), which can only show the phase boundary where spontaneous resolution takes place. It confirms that near the transition, chiral fluctuations become large and hence also the chiral susceptibility that translates to a large helical twisting power. Small differences in the critical exponent for the chiral susceptibility that we find for temperatures near the critical end point and much below that indicate that we cannot rule out that it may be tricritical point. See the SI..
@@
Even not all that close to the transition, the increased susceptibility can be quite significant as Fig. 2 (b) confirms. In the figure, the enantiomeric excess $\eta$ is shown as a function of the dimensionless chiral bias field for different temperatures but fixed chiral interaction strength $K/J=0.333$. In our Monte Carlo simulations, we varied $g/k_{\mathrm{B}}T$ from $0.0$ to $0.25$, $K/J$ from 0 to $0.5$, and $T/T_{IN}$ from $0.668$ to $0.267$, where we annealed $2\times 10^{6}$ Monte Carlo steps at each temperature. In agreement with the theory, we find that the chiral susceptibility goes up with decreasing temperature and with increasing chiral interaction strength $K/J$, as the Fig. S1 shows. According to our theory, both $T$ and $K/J$ enter because $\eta$ not only depends on $g/k_{\mathrm{B}}T$ but also on the value of $\chi$ that itself is a function of $T/T_{\mathrm{IN}}$ and $K/J$.
What Maier-Saupe theory cannot show, and our simulations can, is the spatial structure of the phases and their coarsening behavior. Fig. 4 (a) and (b) show $xz$ cross-sections of a system with size $N=128$, with open boundary conditions on all sides. Both simulations were annealed to a temperature of $T/T_{IN}=0.067$ starting from $T/T_{IN}=1.0$, in $30$ steps of $2\times 10^{6}$ Monte Carlo steps each. In the figures, right-handed ${\eta_{i}}=+1$ spins are shown in red and left-handed ${\eta_{i}}=-1$ spins are shown in blue. Fig. 4(a) shows the racemic nematic phase at low value of $K/J=0.167$, with short range chiral correlations, but no long range order and no induced cholesteric twist. Fig. 4(b) is at a higher value of $K/J=0.5$, showing coexisting layers of left- and right-handed cholesterics approximately two pitch lengths each. A more detailed image can be found in Fig. S3, demonstrating that under these conditions the interface between left- and right-handed cholesteric layers is very sharp. The cholesteric domains form with the pitch axis oriented along a body diagonal of the cubic lattice, see also Fig. 4 (c). This orientation of the pitch axis is allowed by open boundary conditions and is lower in energy than a pitch oriented along the $x$, $y$ or $z$ axes of the cubic simulation volume.
To investigate coarsening behavior in the mirror-image symmetry broken phase, we carried out a simulation of a larger system with $N=480$ and open boundary conditions, for a fixed chiral interaction K/J=0.5{K/J}=0.5 and a temperature $T/T_{IN}=0.067$ that falls in the segregated cholesteric region of the phase diagram. We followed the coarsening over $2\times 10^{6}$ Monte Carlo steps per spin. Our findings are shown in Fig. 4 (c) and (d), where layers with ${\eta_{i}}=1$ are again shown in red and with ${\eta_{i}}=-1$ in blue. Layers with alternating chiral sense form within the first $1\times 10^{5}$ Monte Carlo steps after starting from an initial random configuration as seen in Fig. S5 (a). The sharp interface between left- and right-handed cholesteric layers is apparent even in these early configurations as shown by Fig. S7. These layers widen with time, shown in more detail in the cross sections shown in Fig. S2 and Fig. S5. Initially, multiple domains form with layer normals aligned along different body diagonal axes. Over time, those domains grow, the average layer width increases, and the number of layers drops. Layers annihilate via two mechanisms: (i) motion and annihilation of edge dislocations in the layer structure, and (ii) layer annihilation at corners of the cubic box facilitated by the open boundary conditions. The average layer thickness $L$ we find to increase continuously as $L\sim{t^{\alpha}}$ with scaling exponent $\alpha={0.379}\pm{0.016}$ as shown in Fig. 4 (d). Because layer formation and coarsening are driven by an intrinsically anisotropic surface tension, we do not expect this scaling exponent to be universally valid [37, 39].
As Fig. 4 (c) shows, the domain structure is quite complex, each with layer normals oriented along different body diagonals. Layers formed in early stages of pattern formation have thicknesses less than the cholesteric pitch, increasing to several pitch lengths during coarsening. See Fig. S3, Fig. S6 b), and Fig. S7. Layers less wide than about a pitch do not have a simple, single twisted cholesteric structure but exhibit a double-twist structure. The system remains in a poly-domain structure after $2\times{10^{6}}$ Monte Carlo steps per spin. The dynamics of micro-structural evolution is shown in Supplementary Video S1 and in Fig. S5. Cross-sectional images over the entire simulation length are shown in Fig. S2 with selected insets shown in Fig. S7.
Finally, for values of $K/J=0.833$ and larger, a more complex, disordered double twist geometry without a uniform twist axis emerges, as seen in Fig. S6 (d). The defect-rich micro-structures suggest the formation of blue phases, which have not previously been observed in studies of the chiral Lebwohl-Lasher model. We plan to investigate the details of the blue phase in the uniform chirality Lebwohl-Lasher model in future work.
@@
Discussion and Conclusion
Chiral segregation and deracemization (conversion of a racemate to a mixture with enantiomeric excess) were first observed in racemic mixtures of chiral molecules during crystallization, but can also arise in a variety of other material systems [6]. Indeed, a wide variety of achiral mesogens are known to segregate in chiral domains in different liquid-crystalline phases [33, 44]. For instance, bent-core molecules self-assemble into smectic phases with a chiral arrangement of layer normals, molecular dipole orientations, and molecular tilts, and chemically linked “multimeric” mesogens as well as ferronematic compounds spontaneously segregate in oppositely handed helicoidal twist-bend nematic domains [28, 21, 35, 30, 19].
Current theoretical understanding of why achiral mesogens break mirror-image symmetry is based, on the one hand, on macroscopic elasticity theory [10, 24], and, on the other, on results of computer simulations that focus in particular on the role of particle shape [32, 29, 7]. The role that configurational fluctuations might have that produce degenerate right- and left-handed twisted conformations has attracted much less attention in the theoretical literature [41]. Such chiral fluctuations might become spatially correlated [44], and drive chiral segregation [42, 34]. Wilson et al. have clearly demonstrated the “contagiousness” of chiral conformational states via atomistic molecular simulations of flexible bent-core mesogens [11, 29]. Our combined mean-field theory and computer simulation study demonstrates that this mechanism can play a key role in driving chiral segregation and giving rise to the large helical twisting power typical of chiral dopants.
The mean-field theory and computer simulations that we have presented confirm that favorable homo-chiral interactions between nematogens cause their chiral conformational states to become correlated and promote segregation in oppositely twisted cholesteric domains. That the predicted mirror-image symmetry transition is a continuous one provides an explanation for the large helical twisting power that chiral dopants can have, even if the transition is hidden, e.g., by a transition to a crystal state (see the SI).
Our Maier-Saupe theory predicts $T_{\mathrm{c}}/T_{\mathrm{IN}}=9.08\hskip 2.84544ptq_{0}^{2}\hskip 2.84544ptS^{2}_{\mathrm{c}}$ for the chiral segregation temperature. Hence, to observe the chiral segregation in the nematic phase, we need this ratio to be close enough to unity. This implies that the pitch of the pure P enantiomer must be $P_{0}=2\pi/q_{0}\approx 10-20$ in units of microscopic length. If we take for this the particle length, then this is an unlikely small number with cholesterics in mind. Helicoidal twist-bend nematics, however, are known to have pitches of this order of magnitude, presumably due to their much larger core and hence stronger chiral interaction. Arguably, this explains why chiral segregation is more often seen in nematogens that support this phase.
Finally, our model presumes the P and M configurations of the nematogen to be the most likely configurations, which for $n$-CB certainly is the case. This need not be for other nematogens, which may have as equally stable or more stable configurations achiral ones. For instance, the nematogen DIO exhibits chiral segregation in the nematic phase, but has two chiral and two achiral configuration of about equal energy [50]. All in all, we do not expect explicitly allowing for non-chiral conformers to alter the physics considerably. We intend to investigate this in future work.